OpenStudy (anonymous):

can anyone simplify the expression 4 on the root of 4x^5 times 4 on the root of 4x^3

OpenStudy (anonymous):

\[\sqrt[4]{4x^5\times 4x^3}\]?

OpenStudy (anonymous):

if so answer is \[\sqrt[4]{16x^8}=2x^2\]

OpenStudy (anonymous):

you're awesome